A rectangular area is 5 m longer as it is wide. One side of a square swimming pool will match exactly the width of the area. After the pool is built, the area left will be only 30 m^2. How wide is the area?
Answer by cleomenius(959)

x(x+5) = x^2 + 30
x^2 + 5x = x^2 + 30 We can subtract x^2 from each side.
5x = 30
x = 6
The width is 6.
_____
check
_____
11*5 =66m^2
6^2 = 36M^2; + 30m^2 = 66m^2
cleomenius.